# Impact of Neutrino Flavor Conversions on Neutron Star Merger Dynamics, Ejecta, Nucleosynthesis, and Multi-Messenger Signals

### v1

@@ I Introduction
Next-generation facilities for gravitational waves (GWs) and electromagnetic observations will enable multi-messenger studies of compact object mergers with unprecedented precision and across broader frequency bands Abbott et al. (2016a, 2017a, 2017b, b); Eichler et al. (1989); Radice et al. (2020); Perego et al. (2017); Ruffert and Janka (1999); Just et al. (2016); Cusinato et al. (2021); Martin et al. (2018); Diamond et al. (2024); Vigna-Gómez et al. (2024); Gupta et al. (2024); Niu et al. (2025). The upcoming observations can shape our understanding of some of the most fundamental questions in modern physics and astrophysics, like the nature of dense nuclear matter in stars Lattimer and Prakash (2004); Du et al. (2022); Burgio et al. (2021); Huth et al. (2022); Huang et al. (2022); Chatziioannou et al. (2024); Ecker et al. (2025); Most et al. (2025); Hammond et al. (2025) and rapid neutron-capture (r-process) nucleosynthesis Lattimer et al. (1977); Meyer (1989); Qian and Woosley (1996); Hoffman et al. (1997); Freiburghaus et al. (1999); Metzger et al. (2010); Goriely et al. (2011); Just et al. (2015); Wu et al. (2016); Drout et al. (2017); Balantekin et al. (2024); Chen et al. (2025). One of the most fascinating events to expect are binary neutron star (BNS) mergers, where the majority of r-process might take place. Currently, the exact compositions and yields of the mass ejection of merger remnants remain uncertain Freiburghaus et al. (1999); Korobkin et al. (2012); Wanajo et al. (2014); Thielemann et al. (2017); Radice et al. (2018); Shibata and Hotokezaka (2019); Ekanger et al. (2023); Ricigliano et al. (2024); Chen et al. (2024); Ma et al. (2025); Jacobi et al. (2025) both observationally and theoretically due to the complexity of the interplay between the merger dynamics, nuclear physics, atomic physics, and neutrino interactions Espino et al. (2024a, b, 2023); Pajkos and Most (2025a); Ng et al. (2024); Foucart et al. (2024); Gieg et al. (2025). Theoretical modeling is essential to quantify these uncertainties and to interpret observations, requiring numerical solutions of Einstein’s equations of general relativity with matter.
@@
Early studies identified the collective neutrino flavor conversions in dense neutrino gases, see also Duan et al. (2010); Capozzi and Saviano (2022); Volpe (2024); Johns et al. (2025) for reviews. In particular, the fast flavor instabilities (FFI), a specific class of collective oscillations, is usually studied using idealized radiation-only models Sawyer (2005); Richers et al. (2021); Xiong et al. (2023); Fiorillo and Raffelt (2024a); Xiong et al. (2025a); Kost et al. (2024); Kneller et al. (2024); Lacroix et al. (2024); Xiong et al. (2025a); Nagakura et al. (2025a); Fiorillo and Raffelt (2025a); Johns and Kost (2025); Laraib and Richers (2025); Fiorillo and Raffelt (2025b, c). Some more recent works start to extend these analyses to static snapshots from classical simulations Wu et al. (2017); Richers (2022); Grohs et al. (2023); Froustey et al. (2024); Grohs et al. (2024); Nagakura et al. (2025b); Fiorillo et al. (2025), investigating both the conditions under which FFIs can develop and approaches to modeling them in classical simulations using quantum moment methods Strack and Burrows (2005); Myers et al. (2022); Kneller et al. (2024); Froustey et al. (2025); Grohs et al. (2025). So far, theoretical studies have shown that flavor instabilities driven by neutrino self-interactions can lead to rapid flavor conversion on nanosecond timescales—far shorter than hydrodynamic timescales—making them nearly impossible to model in dynamical simulations, though in-situ flavor transformation could prevent the development of instabilities with timescales as short as those predicted by post-processing models using classical transport Nagakura and Zaizen (2022, 2023); Fiorillo and Raffelt (2024a); Liu et al. (2025a); Fiorillo and Raffelt (2024b); Xiong et al. (2024); Johns et al. (2025).
@@
Despite the challenges of modeling flavor conversions in global simulations, several studies have explored their potential impact in post-merger accretion disks Li and Siegel (2021); Just et al. (2022); Fernández et al. (2022); Mukhopadhyay et al. (2024), including a recent work Lund et al. (2025) that incorporates fast flavor conversions in an angle-dependent neutrino transport scheme. These works have found that, depending on the lifetime of the remnant neutron star Fernández et al. (2022), neutrino flavor conversions can make the disk winds more neutron rich, thus enhancing the r-process in the outflow. However, they focus on a phase of the evolution that does not encompass the highly dynamical merger phase—when the structure and composition of the disk are initially set Camilletti et al. (2024). In-situ parameterized flavor transformation has also been realized in the context of core-collapse supernovae, with results suggesting its important impacts on dynamics and nucleosynthesis Xiong et al. (2020); Ehring et al. (2023a, b); Xiong et al. (2025a); Wang and Burrows (2025); Akaho et al. (2025); Padilla-Gay et al. (2025); Bhattacharyya et al. (2025); Xiong et al. (2025b); Cornelius et al. (2025). Furthermore, there is growing interest in comparing the occurrence and growth rates of FFI and other relevant mechanisms Padilla-Gay et al. (2025), such as collisional flavor instabilities Nagakura et al. (2025b); Wang et al. (2025). Overall, the exact neutrino flavor evolution triggered by a combination of various possible instabilities with different characteristic times remains elusive. Efforts need to be made in both dynamical timescale with in-situ simulations and sub-nano second timescale with quantum kinetics to understand more quantitatively the outcome of the neutrino flavor conversions in dense environments.
@@ II.1 Simulation setup
Our initial configurations consist of two equal-mass binaries of non-rotating NSs, each with a gravitational mass of $1.35\,M_{\odot}$. The stars are placed at an initial separation of $45\,\mathrm{km}$. These initial data are generated using the Lorene pseudo-spectral code Gourgoulhon et al. (2001). We employ two EoSs for the two considered binary systems: the DD2 EoS Typel et al. (2010); Hempel and Schaffner-Bielich (2010) and the SFHo EoS Steiner et al. (2013). Adaptive mesh refinement (AMR) Berger and Oliger (1984); Berger and Colella (1989) is used by employing the Carpet driver Schnetter et al. (2004); Reisswig et al. (2013) of the Einstein Toolkit Loffler et al. (2012); Haas et al. (2024), with 7 refinement levels. The innermost refinement region tracks movements of each neutron star prior to merger and later resolves the central region of the remnant. We consider two resolutions, low resolution (LR) and standard resolution (SR), with corresponding finest grid spacings $0.167\,M_{\odot}\simeq 246\,\mathrm{m}$ and $0.125\,M_{\odot}\simeq 184\,\mathrm{m}$.
@@
Simulation Resolution EoS
$\rho_{\mathrm{mixing}}$
$\mathrm{g/cm}^{3}]$
$\tau_{a}$ [ns] Prescription
$t_{\mathrm{coll}}-t_{\mathrm{merg}}$
$\mathrm{ms}]$
$M^{\mathrm{ej}}_{\mathrm{total}}$
$10^{-2}M_{\odot}]$
$M^{\mathrm{ej}}_{Y_{e}<0.25}/M^{\mathrm{ej}}_{\mathrm{total}}$
$M^{\mathrm{ej}}_{Y_{e}<0.15}/M^{\mathrm{ej}}_{\mathrm{total}}$
$v_{\infty}$ [$c$]
@@ II.2 Neutrino flavor conversion prescriptions
where $\tau_{a}$ is the relaxation time to reach a given asymptotic equilibrium state, $f^{\text{eq}}$, starting from an arbitrary neutrino distribution function, $f$. We apply an operator-splitting approach to incorporate the BGK flavor conversion term, $S_{\mathrm{FC}}$, in our M1 scheme. In the flavor mixing relaxation sub-step, taking place after the source term updates and the feed back to the fluid, the neutrino densities, energies and fluxes are updated according to
@@
As aforementioned, the BGK operator requires inputs of relaxation times and asymptotic states, for which we now introduce two treatments. The first treatment is the same as that in our previous study Qiu et al. (2025). This treatment assumes a piecewise-constant, density dependent relaxation time, $\tau_{a}$, such that $\tau_{a}=0.5$ ns for $\rho<\rho_{\mathrm{mixing}}$, and $\tau_{a}=\infty$ otherwise. We notice that infinite relaxation times corresponds to the absence of flavor conversions. The corresponding asymptotic states, which are motivated by quantum many-body calculations and possible beyond standard model effects Martin et al. (2023), are noted as ‘MB’ (short for many-body), with the assumption that
@@
Figure 3: Post-merger neutrino mean energy for $\nu_{e}$, $\bar{\nu}_{e}$, and $\nu_{x}$, respectively, from left to right panels in each row. The upper row shows results of simulations considering DD2 EoS and the lower row shows those of SFHo EoS. We mark the collapse times for the SFHo runs with dashed lines. For the no-mixing models, heavy-lepton neutrinos generally have higher mean energies, indicating their smaller decoupling radii compared to the electron (anti)neutrinos. For the neutrino mixing models DD2(SFHo)-11, DD2(SFHo)-13 and DD2(SFHo)-FFI, the mean energies of all flavors are comparable, because flavor conversions can exchange energy of different neutrinos. The mean energy of $\nu_{e}$ seems a bit smaller, but it could be due to the fact that high energy $\nu_{e}$ are absorbed more easily by the neutron rich material and the spectrum tends to become a bit softer. The mixing models mostly have higher $\nu_{e}$ and $\bar{\nu}_{e}$ mean energies and lower $\nu_{x}$ mean energy compared to the no-mixing model.
@@ III.1 General dynamics and ejecta properties
We report in Table 1 some of the main ejecta properties of all the simulations. Note that we keep track of the mass ejection and compute the unbound mass based on geodesic criterion, i.e., the matter has to have the time component of its 4-velocity smaller than 1-1. We see that the runs with BH remnants have larger total ejecta mass, and faster ejecta, due to the softer EoS. We also show the ejecta electron fraction profiles and nucleosynthesis yields for both DD2 and SFHo EoSs with 4 models in Fig. 4. In particular, neutrino flavor conversions make the ejecta more neutron rich in the DD2(SFHo)-11, DD2(SFHo)-13, and DD2(SFHo)-FFI simulations compared to the DD2(SFHo)-NM case. Remarkably, the very neutron-rich ejecta fraction ($Y_{e}<0.15$, roughly the threshold for actinides production) can differ by a factor of 3 between the DD2-11 and DD2-NM models, while all the neutron-rich ejecta fraction ($Y_{e}<0.25$, roughly the threshold for lanthanides production) also differs by about a factor of 3. These values likely represent an upper bound on the discrepancy between neutrino-mixing and no-mixing scenarios. In contrast, the same comparison with the SFHo EoS (SFHo-11 vs. SFHo-NM) shows only about a $\sim 100\%$ difference for the $Y_{e}<0.15$ ejecta, and roughly a $\sim 50\%$ difference for the $Y_{e}<0.25$ ejecta. The relative order between DD2(SFHo)-11 and DD2(SFHo)-13 is reversed for the two different EoSs runs. And the amplification of $Y_{e}<0.15$ ejecta in the FFI models is smaller than that of simple density-threshold neutrino mixing models. This is because in the DD2(SFHo)-FFI models, their instability criterion does not flag the entire volume to have neutrino mixing. For the nucleosynthesis yields, the DD2-11 simulation produces about an order of magnitude more lanthanides and heavier elements than the DD2-NM simulation, when the yields are normalized to $A=85$. In the SFHo simulations, the difference is smaller but still reaches up to a factor of five between the SFHo-11 and SFHo-NM relative heavy element yields. Overall, the amount of low-Ye ejecta correlates with the amount of heavy element production.
@@ III.2 Neutrino flavor evolution
We show snapshots on xz-planes of the ELN flux factors (defined in Eq.18) at $\sim$5 ms post-merger for the SFHo-NM, SFHo-11 and SFHo-FFI simulations in Fig. 7. We find that in all the models, there are scattered regions that would be unstable to FFIs. The magnitudes of the ELN flux factors are slightly smaller in the two mixing models, i.e., the SFHo-11 and SFHo-FFI than those in the SFHo-NM. The precise locations of the instabilities differ, likely due to stochasticity across different simulations. However, the overall patterns of the ELN flux factors exhibit qualitatively similar behavior among the three models. The instabilities in polar regions could possibly be explained by the findings in Nagakura et al. (2025b), which discussed correlations between the neutrino chemical potential and the ELN crossing, and suggested that the instabilities could be generated by $\nu_{e}$ contamination. Given that the FFI models explicitly target regions where the ELN flux factor is above 1, one would expect the neutrino flavor mixing to remove such instabilities by pushing them towards flavor equilibrium Zaizen and Nagakura (2023), as the prescriptions for equilibrium states we consider in the BGK model decrease the ELN flux factor. However, our results show that flavor conversions do not necessarily eliminate the instabilities that trigger them. The fact that instabilities remain indicates that the dynamics generating the instabilities must pump the distribution up to unstable levels very rapidly, prevent the system from settling into a stable flavor equilibrium.

### v2

@@ I Introduction
Next-generation facilities for gravitational waves (GWs) and electromagnetic observations will enable multi-messenger studies of compact object mergers with unprecedented precision and across broader frequency bands Abbott and others (2016a, 2017b, 2017a, b); Eichler et al. (1989); Radice et al. (2020); Perego et al. (2017); Ruffert and Janka (1999); Just et al. (2016); Cusinato et al. (2021); Martin et al. (2018); Diamond et al. (2024); Vigna-Gómez and others (2024); Gupta and others (2024); Niu and others (2025). The upcoming observations can shape our understanding of some of the most fundamental questions in modern physics and astrophysics, like the nature of dense nuclear matter in stars Lattimer and Prakash (2004); Du et al. (2022); Burgio et al. (2021); Huth and others (2022); Huang et al. (2022); Chatziioannou et al. (2024); Ecker et al. (2025); Most et al. (2025); Hammond and others (2025) and rapid neutron-capture (r-process) nucleosynthesis Lattimer et al. (1977); Meyer (1989); Qian and Woosley (1996); Hoffman et al. (1997); Freiburghaus et al. (1999); Metzger et al. (2010); Goriely et al. (2011); Just et al. (2015); Wu et al. (2016); Drout and others (2017); Balantekin et al. (2024); Chen et al. (2025). One of the most fascinating events to expect are binary neutron star (BNS) mergers, where the majority of r-process might take place. Currently, the exact compositions and yields of the mass ejection of merger remnants remain uncertain Freiburghaus et al. (1999); Korobkin et al. (2012); Wanajo et al. (2014); Thielemann et al. (2017); Radice et al. (2018); Shibata and Hotokezaka (2019); Ekanger et al. (2023); Ricigliano et al. (2024); Chen et al. (2024); Ma et al. (2025); Jacobi et al. (2025) both observationally and theoretically due to the complexity of the interplay between the merger dynamics, nuclear physics, atomic physics, and neutrino interactions Espino et al. (2024a, b, 2023); Pajkos and Most (2025b); Ng et al. (2024); Foucart et al. (2024); Gieg et al. (2025). Theoretical modeling is essential to quantify these uncertainties and to interpret observations, requiring numerical solutions of Einstein’s equations of general relativity with matter.
@@
Early studies identified the collective neutrino flavor conversions in dense neutrino gases, see also Duan et al. (2010); Capozzi and Saviano (2022); Volpe (2024); Johns et al. (2025) for reviews. In particular, the fast flavor instabilities (FFI), a specific class of collective oscillations, is usually studied using idealized radiation-only models Sawyer (2005); Richers et al. (2021); Xiong et al. (2023); Fiorillo and Raffelt (2024a); Xiong et al. (2025a); Kost et al. (2024); Kneller et al. (2024); Lacroix et al. (2024); Xiong et al. (2025a); Nagakura et al. (2025b); Fiorillo and Raffelt (2025b); Johns and Kost (2025); Laraib and Richers (2025); Fiorillo and Raffelt (2025c, a). Some more recent works start to extend these analyses to static snapshots from classical simulations Wu et al. (2017); Richers (2022); Grohs et al. (2023); Froustey et al. (2024); Grohs et al. (2024); Nagakura et al. (2025a); Fiorillo et al. (2025), investigating both the conditions under which FFIs can develop and approaches to modeling them in classical simulations using quantum moment methods Strack and Burrows (2005); Myers et al. (2022); Kneller et al. (2024); Froustey et al. (2025); Grohs et al. (2025). So far, theoretical studies have shown that flavor instabilities driven by neutrino self-interactions can lead to rapid flavor conversion on nanosecond timescales—far shorter than hydrodynamic timescales—making them nearly impossible to model in dynamical simulations, though in-situ flavor transformation could prevent the development of instabilities with timescales as short as those predicted by post-processing models using classical transport Nagakura and Zaizen (2022, 2023); Fiorillo and Raffelt (2024a); Liu et al. (2025b); Fiorillo and Raffelt (2024b); Xiong et al. (2024); Johns et al. (2025).
@@
Despite the challenges of modeling flavor conversions in global simulations, several studies have explored their potential impact in post-merger accretion disks Li and Siegel (2021); Just et al. (2022); Fernández et al. (2022); Mukhopadhyay et al. (2024), including a recent work Lund et al. (2025) that incorporates fast flavor conversions in an angle-dependent neutrino transport scheme. These works have found that, depending on the lifetime of the remnant neutron star Fernández et al. (2022), neutrino flavor conversions can make the disk winds more neutron rich, thus enhancing the r-process in the outflow. However, they focus on a phase of the evolution that does not encompass the highly dynamical merger phase—when the structure and composition of the disk are initially set Camilletti et al. (2024). In-situ parameterized flavor transformation has also been realized in the context of core-collapse supernovae, with results suggesting its important impacts on dynamics and nucleosynthesis Xiong et al. (2020); Ehring et al. (2023a, b); Xiong et al. (2025a); Wang and Burrows (2025); Akaho et al. (2025); Padilla-Gay et al. (2025); Bhattacharyya et al. (2025); Xiong et al. (2025b); Cornelius et al. (2025). Furthermore, there is growing interest in comparing the occurrence and growth rates of FFI and other relevant mechanisms Padilla-Gay et al. (2025), such as collisional flavor instabilities Nagakura et al. (2025a); Wang et al. (2025). Overall, the exact neutrino flavor evolution triggered by a combination of various possible instabilities with different characteristic times remains elusive. Efforts need to be made in both dynamical timescale with in-situ simulations and sub-nano second timescale with quantum kinetics to understand more quantitatively the outcome of the neutrino flavor conversions in dense environments.
@@ II.1 Simulation setup
Our initial configurations consist of two equal-mass binaries of non-rotating NSs, each with a gravitational mass of $1.35\,M_{\odot}$. The stars are placed at an initial separation of $45\,\mathrm{km}$. These initial data are generated using the Lorene pseudo-spectral code Gourgoulhon et al. (2001). We employ two EoSs for the two considered binary systems: the DD2 EoS Typel et al. (2010); Hempel and Schaffner-Bielich (2010) and the SFHo EoS Steiner et al. (2013). Adaptive mesh refinement (AMR) Berger and Oliger (1984); Berger and Colella (1989) is used by employing the Carpet driver Schnetter et al. (2004); Reisswig et al. (2013) of the Einstein Toolkit Loffler and others (2012); Haas et al. (2024), with 7 refinement levels. The innermost refinement region tracks movements of each neutron star prior to merger and later resolves the central region of the remnant. We consider two resolutions, low resolution (LR) and standard resolution (SR), with corresponding finest grid spacings $0.167\,M_{\odot}\simeq 246\,\mathrm{m}$ and $0.125\,M_{\odot}\simeq 184\,\mathrm{m}$.
@@
Simulation Resolution EoS $\rho_{\mathrm{mixing}}$ $\mathrm{g/cm}^{3}]$ $\tau_{a}$ [ns] Prescription $t_{\mathrm{coll}}-t_{\mathrm{merg}}$ $\mathrm{ms}]$ $M^{\mathrm{ej}}_{\mathrm{total}}$ $10^{-2}M_{\odot}]$ $M^{\mathrm{ej}}_{Y_{e}<0.25}/M^{\mathrm{ej}}_{\mathrm{total}}$ $\%$] $M^{\mathrm{ej}}_{Y_{e}<0.15}/M^{\mathrm{ej}}_{\mathrm{total}}$ $\%$] $v_{\infty}$ [$c$]
@@ II.2 Neutrino flavor conversion prescriptions
where $\tau_{a}$ is the relaxation time to reach a given asymptotic equilibrium state, $f^{\text{eq}}$, starting from an arbitrary neutrino distribution function, $f$. In particular, the BGK subgrid model provides a more accurate treatment than the effective, instantaneous flavor equilibration prescriptions used in, e.g., Li and Siegel (2021); Just et al. (2022), since those methods would introduce errors by approximating continuous evolution as a sequence of discrete instabilities Urquilla and Johns (2025).
We apply an operator-splitting approach to incorporate the BGK flavor conversion term, $S_{\mathrm{FC}}$, in our M1 scheme. In the flavor mixing relaxation sub-step, taking place after the source term updates and the feed back to the fluid, the neutrino densities, energies and fluxes are updated according to
@@
As aforementioned, the BGK operator requires inputs of relaxation times and asymptotic states, for which we now introduce two treatments. The first treatment is the same as that in our previous study Qiu et al. (2025). This treatment assumes a piecewise-constant, density dependent relaxation time, $\tau_{a}$, such that $\tau_{a}=0.5$ ns for $\rho<\rho_{\mathrm{mixing}}$, and $\tau_{a}=\infty$ otherwise. We notice that infinite relaxation times corresponds to the absence of flavor conversions. The choice of $\tau_{a}=0.5$ ns is motivated by the fact that it is comparable to the typical flavor conversion timescales reported in local linear stability analysis of fast flavor conversions in neutron star mergers Froustey et al. (2024). The corresponding asymptotic states, which are motivated by quantum many-body calculations and possible beyond standard model effects Martin et al. (2023), are noted as ‘MB’ (short for many-body). MB assumes detailed balance among the pairwise reactions, such that the rate—which depends on the product of the reactant densities—for converting one flavor pair into another flavor pair equals the rate for the reversed reaction, so
@@
Figure 3: Post-merger neutrino mean energy for $\nu_{e}$, $\bar{\nu}_{e}$, and $\nu_{x}$, respectively, from left to right panels in each row. The upper row shows results of simulations considering DD2 EoS and the lower row shows those of SFHo EoS. We mark the collapse times for the SFHo runs with dashed lines. For the no-mixing models, heavy-lepton neutrinos generally have higher mean energies, indicating their smaller decoupling radii compared to the electron (anti)neutrinos. For the neutrino mixing models DD2(SFHo)-11, DD2(SFHo)-13 and DD2(SFHo)-FFI, the mean energies of all flavors are comparable, because flavor conversions can exchange energy of different neutrinos. The mean energy of $\nu_{e}$ seems a bit smaller, but it could be due to the fact that high energy $\nu_{e}$ are absorbed more easily by the neutron rich material and the spectrum tends to become a bit softer. The mixing models mostly have higher $\nu_{e}$ and $\bar{\nu}_{e}$ mean energies and lower $\nu_{x}$ mean energy compared to the no-mixing model. Note that the sharp dip in the DD2-FFI model at around $\sim$3.5 ms post-merger is caused by a transient outburst in the $\nu_{x}$ number flux. This outburst is generated when a parcel of material crosses a refinement-level corner in the mesh, producing a short–lived numerical artifact rather than a physical signal.
@@ III.1 General dynamics and ejecta properties
We report in Table 1 some of the main ejecta properties of all the simulations. Note that we keep track of the mass ejection and compute the unbound mass based on geodesic criterion, i.e., the matter has to have the time component of its 4-velocity smaller than 1-1. We see that the runs with BH remnants have larger total ejecta mass, and faster ejecta, due to the softer EoS. We also show the ejecta electron fraction profiles and nucleosynthesis yields for both DD2 and SFHo EoSs with 4 models in Fig. 4. In particular, neutrino flavor conversions make the ejecta more neutron rich in the DD2(SFHo)-11, DD2(SFHo)-13, and DD2(SFHo)-FFI simulations compared to the DD2(SFHo)-NM case. Remarkably, the very neutron-rich ejecta fraction ($Y_{e}<0.15$, roughly the threshold for actinides production) can differ by a factor of 3 between the DD2-11 and DD2-NM models, while all the neutron-rich ejecta fraction ($Y_{e}<0.25$, roughly the threshold for lanthanides production) also differs by about a factor of 3. These values likely represent an upper bound on the discrepancy between neutrino-mixing and no-mixing scenarios. In contrast, the same comparison with the SFHo EoS (SFHo-11 vs. SFHo-NM) shows only about a $\sim 100\%$ difference for the $Y_{e}<0.15$ ejecta, and roughly a $\sim 50\%$ difference for the $Y_{e}<0.25$ ejecta. The relative order between DD2(SFHo)-11 and DD2(SFHo)-13 is reversed for the two different EoSs runs. And the amplification of $Y_{e}<0.15$ ejecta in the FFI models is smaller than that of simple density-threshold neutrino mixing models. This is because in the DD2(SFHo)-FFI models, their instability criterion does not flag the entire volume to have neutrino mixing. For the nucleosynthesis yields, the DD2-11 simulation produces about an order of magnitude more lanthanides and heavier elements than the DD2-NM simulation, when the yields are normalized to $A=85$. In the SFHo simulations, the difference is smaller but still reaches up to a factor of five between the SFHo-11 and SFHo-NM relative heavy element yields. In particular, we find quantitative agreement with Just et al. (2022) regarding the heavy-element yields in the SFHo-FFI model, with the difference between the mixing and no-mixing cases being around 100%, despite our ejecta properties being measured at earlier, closer-to-merger timescales. Overall, the amount of low-Ye ejecta correlates with the amount of heavy element production.
@@ III.2 Neutrino flavor evolution
We show snapshots on xz-planes of the ELN flux factors (defined in Eq.18) at $\sim$5 ms post-merger for the SFHo-NM, SFHo-11 and SFHo-FFI simulations in Fig. 7. We find that in all the models, there are scattered regions that would be unstable to FFIs. The magnitudes of the ELN flux factors are slightly smaller in the two mixing models, i.e., the SFHo-11 and SFHo-FFI than those in the SFHo-NM. The precise locations of the instabilities differ, likely due to stochasticity across different simulations. However, the overall patterns of the ELN flux factors exhibit qualitatively similar behavior among the three models. The instabilities in polar regions could possibly be explained by the findings in Nagakura et al. (2025a), which discussed correlations between the neutrino chemical potential and the ELN crossing, and suggested that the instabilities could be generated by $\nu_{e}$ contamination. Given that the FFI models explicitly target regions where the ELN flux factor is above 1, one would expect the neutrino flavor mixing to remove such instabilities by pushing them towards flavor equilibrium Zaizen and Nagakura (2023), as the prescriptions for equilibrium states we consider in the BGK model decrease the ELN flux factor. However, our results show that flavor conversions do not necessarily eliminate the instabilities that trigger them. The fact that instabilities remain indicates that the dynamics generating the instabilities must pump the distribution up to unstable levels very rapidly, prevent the system from settling into a stable flavor equilibrium.